OpenStudy (ttp):

What is the distance from the coordinate pair one comma negative three to the coordinate pair negative three comma four? Round your answer to the nearest hundredth. three point three one seven point eight one seven point nine three eight point zero six

OpenStudy (ttp):

@texaschic101

OpenStudy (texaschic101):

distance formula : sq rt ((x2 - x1)^2 + (y2 - y1)^2) (1,-3) x1 = 1 and y1 = -3 (-3,4) x2 = -3 and y2 = 4 now we sub distance = sq rt ((-3-1)^2 + (4 - (-3) ^2) distance = sq rt (-4^2 + 7^2) distance = sq rt (16 + 49) distance = sq rt 65 distance = 8.06